# Katherine Atwell*§    Pedram Heydari*†    Anthony Sicilia¶    Malihe Alikhani§

### v3

@@ Abstract.
We introduce a Bayesian probabilistic framework, grounded in behavioral economics and rational decision theory, that explicitly separates sycophancy from rational belief updating. Within this framework, we achieve three objectives: (i) a descriptive metric that measures sycophancy while controlling for rational responses to evidence; (ii) a normative metric that quantifies how sycophancy leads models astray from Bayesian-consistent belief updating; and (iii) the ability to apply both metrics in settings without ground-truth labels. Applying our framework across multiple LLMs and three uncertainty-driven tasks, we find robust evidence of sycophantic belief shifts and show that their impact on rationality depends on whether models systematically over- or under-update their beliefs. Finally, we demonstrate that a post-hoc calibration method and two fine-tuning strategies (SFT and DPO) substantially reduce Bayesian inconsistency, with particularly strong improvements under explicit sycophancy prompting.
@@ 1. Introduction
Building on the theoretical foundation of BASIL, the final component of our work moves from measurement to mitigation. We propose and evaluate three distinct interventions designed to enforce the ”shared norms” of belief formation that sycophancy typically disrupts: calibration and post-Training interventions. Our first intervention addresses the issue of model miscalibration—the tendency for LLMs to express over or under-confidence in their initial priors. We introduce a novel multi-step calibration strategy: first, we apply isotonic regression to align the model’s prior beliefs with ground-truth distributions. Then, rather than simply correcting the prior, we use odds-ratio scaling to propagate this correction through the model’s posterior estimates. This ensures that the model’s update remains ”subjectively rational”—internally consistent with its now-calibrated baseline—even in the absence of ground-truth labels for the posterior itself. Furthermore, we investigate whether models can be actively trained to prioritize logical consistency over user-alignment. We explore two post-training strategies that utilize our normative metric as a supervisory signal: Supervised Fine-Tuning (SFT): Models are trained to output the calculated Bayesian-rational posterior, teaching them to ”show their work” in a way that respects the mathematical relationship between their own stated priors and likelihoods. Direct Preference Optimization (DPO): We employ a label-free preference ranking where the ”preferred” completion is the one that minimizes the distance to the model’s own Bayesian-rational posterior. This rewards the model for resisting the ”sycophancy tax” and maintaining its internal logical standard regardless of the user’s expressed opinion.
@@
We apply our Bayesian framework across three tasks involving inherent uncertainty: conversation forecasting, morality judgments, and cultural acceptability judgments (A). We test the following hypotheses:
@@ 2.2. Bayesian Reasoning
Recent literature has found that LLMs can output calibrated probability estimates when prompted to verbalize their estimates to the user (Lin et al., 2022). Indeed, there is evidence that verbalized probabilities are more calibrated than conditional token probabilities. Hence, in this work, we experiment with eliciting verbal probability estimates in the form of a percentage (similar to (Xiong et al., 2023)), which we refer to as direct probing. Taking inspiration from the self-random sampling approach used by (Xiong et al., 2023), we also experiment with what we deem a hybrid approach, where rather than ask the model the same question multiple times, we instead ask it “If we were to ask you 10 times, how many times would you say that the following is true?” This approach is more efficient than self-random sampling, as it does not require the same question to be asked multiple times.
@@ 3.2.2. Quantifying sycophancy
Using these metrics, we quantify the effects of the three transitions shown above (third-party, from Abstract to Third-party belief, user, from Third-party belief to User belief, and total, from Abstract to Third-party belief) on LLMs’ stated beliefs and the rationality of these beliefs.
@@
We also provide results for rate of change in Appendix E, in addition to our formula for calculating rate of change.
@@
We also report the KL divergence between the Bayesian-rational posterior and predicted posterior in Appendix E, which also includes our equation for calculating KL divergence.
@@ 4.2. Eliciting Probability Estimates from LLMs
Brier Score Brier Skill Score
Raw Cal. Raw Cal.
Direct Probing 0.2964 0.1579 -0.0774 0.2840
Hybrid 0.4407 0.2022 -0.8345 0.1366
Direct Probing Samples=5 0.2674 0.1888 -0.2674 0.0910
Table 3. Sycophancy scores using our descriptive measure of sycophancy, detailing the change in probability estimates between the baseline and the intervention designed to probe for sycophancy for the raw and calibrated probabilities (each baseline is described in §3). A higher score indicates more sycophancy, and negative scores indicate a change in the opposite direction from the user’s stated or implied beliefs.
Raw Calibrated
Total Third-Party User Total Third-Party User
@@ 4.3. Baselines
Table 4. Sycophancy scores and Bayesian error for each baseline for the direct probing probability elicitation method (top), and the direct probing with multiple sampled method (bottom) based on the models’ raw and calibrated probability estimates. *denotes statistical significance at $p<0.1$, and **denotes statistical significance at $p<0.05$, using the Wilcoxon Signed Rank Test. Lighter colors represent negative values (indicating a reduction in error due to sycophancy), while darker colors represent larger, positive values (indicating an increase in error due to sycophancy). We find that models consistently demonstrate sycophantic behavior when presented with the user’s beliefs, but that the impacts of sycophancy on Bayesian error are dependent upon the nature of the model’s updates. While models that over-update their beliefs are more likely to have more errors due to sycophancy, this is not the case for models that under-update their beliefs.
@@ 4.4. Calibrating LLMs’ beliefs
To calibrate our priors, we apply isotonic regression on LLMs’ verbalized probability estimates using the ground-truth labels for each outcome. Table 2 shows a reduction in calibration error across all of our techniques for eliciting probability when this method is applied.
@@ 4.5. Post-training to reward rational behavior in LLMs
We experiment with two different post-training approaches, both of which utilize our metric to reward Bayesian rationality in models: a supervised finetuning approach (SFT) and direct preference optimization (DPO) approach. We first experiment with finetuning models to output Bayesian-rational posteriors based on their elicited priors, likelihoods, and evidence using SFT. Lin et al. (2022) find the models can be finetuned to verbalize more well-calibrated probabilities, and we wish to investigate whether a similar approach can yield more Bayesian-rational models. Next, we experiment with DPO, where in lieu of using human preference labels, we rank generated answers based on their consistency with their previously-expressed beliefs.
@@ 5. Results
5.1. Stating the user’s belief in a given outcome will significantly shift LLMs’ stated beliefs towards this outcome, compared to when no outside beliefs are provided and when a third-party belief is provided.
To test this hypothesis, we use our descriptive sycophancy metric to capture belief shifts between the Abstract and Third-Party Belief case ($\Delta_{third}$), and between the Abstract and User-beliefs case ($\Delta_{total}$) (see §3.2.1 for details. Prior research has demonstrated significant shifts in models’ stated beliefs for the latter case, and we hypothesize that significant shifts will also occur in the former case. As shown in Table 3, we find that, overall, stating the user’s belief significantly changes models’ beliefs towards the user’s stated beliefs when compared to both the abstract case ($LOC_{total}$) and the Third-party belief case ($LOC_{user}$). This result supports our above hypothesis that both settings will elicit belief changes in models.
@@
In Table 4, we report the sycophancy scores (Rate of Change and Log Odds Change) and the corresponding shifts in Bayesian error ($\Delta$ KL Divergence and $\Delta$ RMSE) for all baselines when transitioning from the Abstract to the Sycophancy condition. Across the aggregate results (“All”), we observe significant increases in RMSE for several baselines and no overall significant decreases. However, the data reveals a critical nuance when we disaggregate by updating style. Consistent with our hypothesis, we observe significant increases in Bayesian error in instances where models over-update their beliefs—i.e., when the predicted posterior already exceeds the Bayesian-rational posterior ($\hat{P}(X|E)>P^{*}(X|E)$). This trend holds across the majority of our studied baselines. Conversely, we observe consistent decreases in RMSE during under-updating scenarios ($\hat{P}(X|E)<P^{*}(X|E)$), with only one exception. We characterize this latter effect not as a functional improvement in reasoning, but as a compensatory distortion: the social pressure of sycophancy pushes an otherwise ”stubborn” or conservative model toward the rational posterior by accident. These results validate our hypothesis that the normative impact of sycophancy is directionally dependent, acting as an additional source of error for over-confident models while masking underlying reasoning deficits in under-confident ones.
@@
Figure 3. Root mean square error between Bayesian-rational posterior and predicted posterior for raw probabilities, calibrated priors only, and our calibration technique with scaled posteriors. While only calibrating priors increases Bayesian error, our technique of calibrating priors and scaling the posteriors reduces Bayesian error for both the base and sycophancy cases.
5.4. Hypothesis IV: Post training that directly rewards Bayesian consistent updates reduces Bayesian inconsistency overall and reduces the extra inconsistency caused by sycophancy
As shown in Table 4, we find that our both our SFT and DPO baselines reduce the gap in RMSE between the Abstract and Sycophancy cases, and Figure 4 shows a substantial reduction in error for our SFT baseline across the Abstract, Third-Party Belief, and Sycophancy cases, and a similar (but slightly lower) reduction in error for the DPO baseline. This is particularly true for the Third-Party Belief and Sycophancy cases, suggesting that our approaches can mitigate some of the negative impacts of sycophancy on Bayesian rationality, beyond simply making models more Bayesian-rational overall.
Figure 4. Root mean square error between Bayesian-rational posterior and predicted posterior for our base Llama 3.2:1b baseline and our finetuned baseline using SFT. We find that, for both the base and sycophancy cases, our metric reduces Bayesian error by a substantial margin.
@@ 6. Conclusions and Future Work
In this work, we introduced BASIL, a Bayesian framework designed to disentangle sycophantic behavior from rational belief updating in LLMs. By adopting a two-dimensional approach, we quantified both the descriptive magnitude of belief shifts and the normative impact of these shifts on a model’s internal logical consistency. Our results confirm that direct sycophancy probing significantly distorts a model’s stated posterior. Crucially, we demonstrate that the normative impact of sycophancy is directionally dependent: while sycophancy consistently increases Bayesian error in over-updating models, it can act as a compensatory distortion in under-updating models, masking underlying reasoning flaws by pushing the model toward a ”rational” posterior for the wrong reasons.
Most significantly, we identified two robust pathways for reducing Bayesian error. First, our novel calibration strategy—propagating calibrated priors through the posterior via odds-ratio scaling—effectively reduces error in both baseline and sycophantic contexts. This method remains functional even when ground-truth labels are only available for the priors. Second, we developed label-free post-training interventions. By utilizing supervised fine-tuning (SFT) and Direct Preference Optimization (DPO) to reward internal Bayesian consistency rather than human-labeled preferences, we significantly reduced reasoning errors and mitigated the ”extra” inconsistency caused by social pressure.
@@ D.3.2. Hybrid
Appendix E Full Results: Sycophancy
E.1. KL Divergence

### v4

@@ Abstract.
We introduce a Bayesian probabilistic framework, grounded in behavioral economics and rational decision theory, that explicitly separates sycophancy from rational belief updating. Within this framework, we propose two group-truth-independent metrics for studying sycophancy: (i) a descriptive metric that measures sycophancy while controlling for rational responses to evidence, and (ii) a normative metric that quantifies how sycophancy leads models astray from Bayesian-consistent belief updating. Applying our framework across multiple LLMs and three uncertainty-driven tasks, we find robust evidence of sycophantic belief shifts and show that their impact on rationality depends on whether models systematically over- or under-update their beliefs, with most baselines demonstrating significant increases in error due to sycophancy when the model over-updates. Finally, we propose a novel post-hoc calibration method and two fine-tuning strategies that reward Bayesian-rational updating (BayesSFT and BayesDPO). We find evidence that post-hoc calibration significantly reduces Bayesian error, and observe significant reductions in both sycophancy and Bayesian error associated with our novel fine-tuning methods.
@@ 1. Introduction
Building on the theoretical foundation of BASIL, the final component of our work moves from measurement to mitigation. We propose and evaluate three distinct interventions designed to enforce the ”shared norms” of belief formation that sycophancy typically disrupts: calibration and two post-training interventions. Our first intervention addresses the issue of model miscalibration—the tendency for LLMs to express over or under-confidence in their initial priors. We introduce a novel multi-step calibration strategy: first, we apply isotonic regression to align the model’s prior beliefs with ground-truth distributions. Then, rather than simply correcting the prior, we use odds-ratio scaling to propagate this correction through the model’s posterior estimates. This ensures that the model’s update remains ”subjectively rational”—internally consistent with its now-calibrated baseline—even in the absence of ground-truth labels for the posterior itself. Furthermore, we investigate whether models can be actively trained to prioritize logical consistency over user-alignment. We propose two post-training strategies that utilize our normative metric as a supervisory signal: BayesSFT, where models are trained to output predictions consistent with their base beliefs using supervised finetuning (SFT), and BayesDPO, a modification of direct preference optimization (DPO) with a label-free preference ranking where the “preferred” completion is the one that minimizes the distance to the model’s own Bayesian-rational posterior. This rewards the model for resisting the ”sycophancy tax” and maintaining its internal logical standard regardless of the user’s expressed opinion.
@@
We apply our Bayesian framework across three tasks involving inherent uncertainty: conversation forecasting, morality judgments, and cultural acceptability judgments (§4.1). We test the following hypotheses:
@@ 2.2. Bayesian Reasoning
Recent literature has found that LLMs can output calibrated probability estimates when prompted to verbalize their estimates to the user (Lin et al., 2022). Indeed, there is evidence that verbalized probabilities are more calibrated than conditional token probabilities. Hence, in this work, we experiment with eliciting verbal probability estimates in the form of a percentage (similar to (Xiong et al., 2023)), which we refer to as direct probing. Here, we set the temperature at zero, in order to obtain the model’s probability estimate under greedy sampling. Taking inspiration from the self-random sampling approach used by (Xiong et al., 2023), we also experiment with what we deem a hybrid approach, where rather than ask the model the same question multiple times, we instead ask it “If we were to ask you 10 times, how many times would you say that the following is true?” This approach is more efficient than self-random sampling, as it does not require the same question to be asked multiple times. Here, we also set the temperature to 0. Finally, we experiment with a combination of direct probing and self-random sampling, where we directly probe a model multiple times, setting the temperature above zero to ensure stochasticity. We refer to this approach as direct probing with multiple samples.
@@ 3.2.2. Quantifying sycophancy
Using these metrics, we quantify the effects of the three transitions shown above (third-party, from Abstract to Third-party belief, user, from Third-party belief to User belief, and total, from Abstract to User belief) on LLMs’ stated beliefs and the rationality of these beliefs.
@@
We also provide results for rate of change in Appendix F, in addition to our formula for calculating rate of change.
@@
We also report the KL divergence between the Bayesian-rational posterior and predicted posterior in Appendix F, which also includes our equation for calculating KL divergence.
@@ 4.2. Eliciting Probability Estimates from LLMs
Table 2. Bayesian error (RMSE) for all pretrained baselines across all three confidence elicitation methods (direct probing, hybrid, direct probing with multiple samples). The hybrid method is associated with the most Bayesian error on average, and model size does not appear to have much impact on error.
Direct Probing Hybrid Direct probing (samples=5)
Abstract Third-p. belief User belief Abstract Third-p. belief User belief Abstract Third-p. belief User belief
[…]
@@
Table 3. Sycophancy scores for the direct probing (left) and direction probing with multiple samples (right) elicitation methods using our descriptive measure of sycophancy, detailing the change in probability estimates between the baseline and the intervention designed to probe for sycophancy for the raw and calibrated probabilities (each baseline is described in §3). A higher score (darker) indicates more sycophancy, and negative scores (light) indicate a change in the opposite direction from the user’s stated or implied beliefs. * denotes statistical significance at $p<0.1$, and ** denotes statistical significance at $p<0.05$, using the Wilcoxon Signed Rank Test. llama3.2:1b+SFT and llama3.2:1b+DPO refer to our llama3.2:1b baseline, post-trained using our BayesSFT and BayesDPO method, respectively.
Direct Probing Direct Probing (Multiple Samples=5)
Raw Calibrated Raw Calibrated
Total 3rd-P User Total 3rd-P User Total 3rd-P User Total 3rd-P User
@@ 4.3. Baselines
Table 4. Change in Bayesian error due to sycophancy for each baseline for the direct probing probability elicitation method (top), and the direct probing with multiple samples method (bottom) based on the models’ raw and calibrated probability estimates. * denotes statistical significance at $p<0.1$, and ** denotes statistical significance at $p<0.05$, using the Wilcoxon Signed Rank Test. Bayesian error is calculated as the root mean squared error between the predicted posterior and the Bayesian-rational posterior. Lighter colors represent smaller values, with the lightest being negative (indicating a reduction in error due to sycophancy), while darker colors represent larger, positive values (indicating an increase in error due to sycophancy). We find that models consistently demonstrate sycophantic behavior when presented with the user’s beliefs, but that the impacts of sycophancy on Bayesian error are dependent upon the nature of the model’s updates (consistent with hypothesis II).
@@ 4.4. Calibrating LLMs’ beliefs
To calibrate our priors, we apply isotonic regression on LLMs’ verbalized probability estimates using the ground-truth labels for each outcome. Table 6 in Appendix E shows reduced calibration error across all of our techniques for eliciting probability when this method is applied.
@@ 4.5. Post-training to reward rational behavior in LLMs
We experiment with two different post-training approaches, both of which utilize our metric to reward Bayesian rationality in models: a supervised finetuning approach, which we call BayesSFT, and a modified direct preference optimization approach, which we call BayesDPO. We describe these approaches below:
  • BayesSFT: Motivated by Lin et al. (2022), who have found that LLM can be finetuned to verbalize more well-calibrated probabilities, we experiment with finetuning our models on their most Bayesian-rational predicted posteriors. Based on their initial predicted priors, posteriors, likelihoods, and alternative likelihoods under greedy sampling, we obtain the 200 data points with the most Bayesian-rational predicted posteriors for each dataset in the Abstract setting (600 total). Using this data, we finetune our model to output predicted posteriors for the Abstract, Third-party belief, and User belief cases.
  • BayesDPO: Direct preference optimization (DPO) (Rafailov et al., 2023) allows models to be directly tuned on preference data without the need to train a separate reward model. Here, we propose a modified DPO approach where, instead of using user preference labels, we rank candidate responses based on how Bayesian they are. Based on the model’s initial predicted priors, posteriors, likelihoods, and alternative likelihoods under greedy sampling, we present the model with two candidate posteriors at each step, with the more Bayesian-rational posterior marked as “chosen”. Our goal is to train a reward model that penalizes inconsistent belief updating.
@@ 5. Results
5.1. Our hybrid method for eliciting model beliefs is associated with the most Bayesian error, and model size appears not to have much impact on Bayesian rationality.
In Table 2, we show the Bayesian error for each baseline’s raw probability estimates, for each probability elicitation method (direct probing, hybrid, and direct probing with multiple samples) and each test case for sycophancy probing (Abstract, Third-party belief, and User belief). We observe that our hybrid probability elicitation method is associated with much more Bayesian error, on average, than the two direct probing methods. In essence, our hybrid method studies how models behave when trying to predict their own behavior. Our results indicate that this is associated with much less belief consistency on average. Future work could compare the errors associated with our hybrid method to those observed for self-random sampling, to compare observed model behavior (self-random sampling) with the model’s own predictions about its behavior (hybrid method). Because our hybrid method is associated with such low belief consistency in general, we focus on the other two methods for the remainder of the paper, with the hybrid results shown in full in Appendix F.
We observe some differences in Bayesian error between very large closed models (gpt-40-mini and claude-haiku-4-5) and very small open-source models (llama 3.2:1b and llama 3.2:3b), with the larger closed models generally exhibiting less error. However, we observe that phi-4-14b achieves comparable results to claude-haiku-4-5, indicating that other factors beyond model size may impact Bayesian rationality in LLMs. Further, we observe that Mistral 7b is associated with the most Bayesian error overall, with more error on average than the two smaller Llama models.
5.2. Hypothesis I: Stating the user’s belief in a given outcome will significantly shift LLMs’ stated beliefs towards this outcome, compared to when no outside beliefs are provided and when a third-party belief is provided.
To test this hypothesis, we use our descriptive sycophancy metric to capture belief shifts between the Abstract and Third-Party Belief case ($\Delta_{third}$), and between the Abstract and User-beliefs case ($\Delta_{total}$) (see §3.2.1 for details. Prior research has demonstrated significant shifts in models’ stated beliefs for the latter case, and we hypothesize that significant shifts will also occur in the former case. As shown in Table 3, we find that, overall, stating the user’s belief significantly changes models’ beliefs towards the user’s stated beliefs when compared to both the abstract case ($LOC_{total}$) and the Third-party belief case ($LOC_{user}$). This result supplements existing literature showing evidence of model sycophancy, while also providing definitive proof that the user’s stated beliefs have an outsized impact on model predictions, even when controlling for the information gain that occurs when a third-party’s beliefs are provided.
@@
In Table 4, we report the shifts in Bayesian error for all baselines when transitioning from the Abstract to the Sycophancy condition ($\Delta_{total}$ RMSE). The aggregate results (“All”) show some significant increases in RMSE, but also some significant decreases. However, the data reveals a critical nuance when we disaggregate by updating style. Consistent with our hypothesis, we observe significant increases in Bayesian error in instances where models over-update their beliefs (when the predicted posterior already exceeds the Bayesian-rational posterior ($\hat{P}(X|E)>P^{*}(X|E)$)) and sycophancy is observed ($\hat{P}^{+S}(X|E)>\hat{P}(X|E)$). This trend holds across the majority of our studied baselines. Conversely, we observe consistent decreases in RMSE during under-updating scenarios ($\hat{P}(X|E)<P^{*}(X|E)$) where sycophancy occurs ($\hat{P}^{+S}(X|E)>\hat{P}(X|E)$) with only one exception. We characterize this latter effect not as a functional improvement in reasoning, but as a compensatory distortion: the social pressure of sycophancy pushes an otherwise ”stubborn” or conservative model toward the rational posterior by accident. These results validate our hypothesis that the normative impact of sycophancy is directionally dependent, acting as an additional source of error for over-confident models while masking underlying reasoning deficits in under-confident ones.
@@
Figure 3. Mean squared error between Bayesian-rational posterior and predicted posterior for raw probabilities, calibrated priors only, and our calibration technique with scaled posteriors, averaged across all pretrained baselines for the direct probing with multiple samples strategy. We report mean squared error here in order to directly calculate confidence intervals (shown as error bars). While only calibrating priors increases Bayesian error, our technique of calibrating priors and scaling the posteriors significantly reduces Bayesian error for both the base and sycophancy cases.
5.5. Hypothesis IV: Post training that directly rewards Bayesian consistent updates reduces sycophancy and Bayesian inconsistency
As shown in Figure 4, we find that our both our BayesSFT and BayesDPO approaches significantly reduce Bayesian error across the Abstract, Third-Party Belief, and Sycophancy cases. This validates the use of these two novel approaches for reducing Bayesian error, and aligns with our prediction that these post-training methods will be associated with less Bayesian inconsistency.
Further, we observe that BayesSFT is associated with a reduction in total sycophancy for the direct probing elicitation method and BayesDPO is associated with a reduction in total sycophancy overall (Table 3). This also holds true when measuring sycophancy for the User case (from Third-party beliefs to User beliefs). This behavior is expected and aligns with our hypothesis; because these two baselines involved tuning on the same predicted posteriors for the Abstract, Third-party belief, and User belief cases, our approach awards consistent probability estimates for each of these three cases.
Figure 4. Mean squared error between Bayesian-rational posterior and predicted posterior for our base Llama 3.2:1b baseline and our finetuned baselines using BayesSFT and BayesDPO, respectively for the direct probing with multiple samples elicitation method. We report mean squared error here in order to directly calculate confidence intervals (shown as error bars). We find that, for both the base and sycophancy cases, both BayesSFT and BayesDPO significantly reduce Bayesian error.
@@ 6. Conclusions and Future Work
In this work, we introduce BASIL, a Bayesian framework designed to disentangle sycophantic behavior from rational belief updating in LLMs. Our framework’s two-dimensional approach quantifies both the descriptive magnitude of belief shifts and the normative impact of these shifts on a model’s internal logical consistency. Our results confirm that direct sycophancy probing significantly distorts a model’s stated posterior, and that including the user’s belief in particular yields a strong shift in the model’s posterior, compared to when no beliefs or a third-party’s beliefs are included. Crucially, we demonstrate that the normative impact of sycophancy is directionally dependent: while sycophancy consistently increases Bayesian error in over-updating models, it can act as a compensatory distortion in under-updating models, masking underlying reasoning flaws by pushing the model toward a ”rational” posterior for the wrong reasons.
Most significantly, we identified two robust pathways for reducing Bayesian error. First, our novel calibration strategy—propagating calibrated priors through the posterior via odds-ratio scaling—effectively reduces error in both baseline and sycophantic contexts. This method remains functional even when ground-truth labels are only available for the priors. Second, we introduce two label-free post-training interventions for improving Bayesian-rationality and reducing the impact of sycophancy. Using our novel approaches, BayesSFT and BayesDPO, to reward internal Bayesian consistency rather than human-labeled preferences, we significantly reduce reasoning errors and reduce shifts in behavior when provided with information about users’ beliefs.
@@ D.3.2. Hybrid
Appendix E Calibration Error: Raw vs. Calibrated Priors
Brier Score Brier Skill Score
Raw Cal. Raw Cal.
Direct Probing 0.2964 0.1579 -0.0774 0.2840
Hybrid 0.4407 0.2022 -0.8345 0.1366
Direct Probing Samples=5 0.2674 0.1888 -0.2674 0.0910
Appendix F Full Results: Sycophancy
F.1. KL Divergence